# Measuring gravitational lensing time delays with quantum information processing

### v1

@@ II.1 Gravitational lensing time delays
When light emitted by a distant source passes near a large distribution of mass on its way to Earth, the path of the light is altered in an effect known as gravitational lensing. For sufficiently large masses, the distortion is large enough to form multiple images of the same source on the sky. However, at lower masses, these images cannot be individually resolved—they overlap with each other and with the true position of the source, causing the source to appear brighter. This transient magnification of a source is known as gravitational microlensing Paczynski (1986). Such microlensing events provide one of the few ways to detect non-luminous astrophysical bodies.
For the purposes of this paper, we will restrict ourselves to a fiducial example, namely a star in the Galactic Bulge of the Milky Way as the source and a dark, isolated object such as a black hole as the lens. This example is of particular interest, as the Galactic Bulge has been the target of decades of ground-based microlensing searches Udalski et al. (2015); Nunota et al. (2025); Park et al. (2018), and such surveys have yielded significant discoveries of dark objects such as isolated black holes Lam et al. (2022); Sahu et al. (2022); Kaczmarek et al. (2025) and free-floating planets Sumi et al. (2011); Mróz et al. (2017, 2018, 2019, 2020, 2020); Kim et al. (2021); Mroz and Poleski (2023); Mróz et al. (2024); Koshimoto et al. (2023); Gould et al. (2022, 2023). In the next few years, future missions such as NASA’s Nancy Grace Roman Space Telescope (henceforth, Roman) Akeson et al. (2019); Penny et al. (2019) and the Chinese National Space Agency’s Earth 2.0 satellite Ge et al. (2022) will conduct the first ever dedicated space-based microlensing surveys. They are expected to discover orders of magnitude more dark astrophysical bodies than current ground-based surveys Penny et al. (2019); Johnson et al. (2020).
However, as powerful a tool as microlensing is for discovering dark astrophysical objects, it suffers from inherent degeneracies that make measuring the underlying properties of the lens, such as its mass, quite challenging Lee (2017). This can be seen from the fact that the primary observable associated with microlensing is a quantity called the Einstein crossing time that corresponds to the approximate duration of the lensing event. The Einstein crossing time is defined as the time for the source to cross the angular Einstein radius $\theta_{E}$, which is the region of the sky surrounding the lens in which the source is magnified by the lens. The Einstein crossing time $t_{E}$ therefore depends on the distance to the lens ($D_{L}$), the distance to the source ($D_{S}$), the mass of the lens ($M$), and the relative proper motion of the lens and source $\mu_{\text{rel}}$:
@@
The impact parameter can be independently measured from the magnification curve, as the magnification in the point-source point-lens regime is Refsdal (1964)
@@
where $u$ varies with time as the source location approaches its minimum impact parameter $u_{0}$ and then moves away from the lens axis. We see that if $u_{0}=0$, the magnification is infinite, as expected Paczynski (1986). Given the magnification curve and the time delay, it is clear that one can solve for $M$, the lens mass, directly, breaking the inherent degeneracy present in the light curve alone.
The behavior of Eqs. (7) and (8), as plotted in Fig. 2, constrains the regime in which microlensing is useful. For a microlensing event to be detectable, we normally require $u\approx 1$ (though with future space-based missions, this may be relaxed). Additionally, note that Eq. (8) is the total magnification, i.e., the sum of the major and minor image contributions given by $A_{\pm}(u)=\frac{1}{2}A(u)\pm\frac{1}{2}$. The analysis is simpler when there is comparable flux from the two images, i.e., $u\approx 1$. For both of these reasons, we mainly consider examples in which $u\approx 1$, hence $\Delta t\approx\frac{8GM}{c^{3}}$ and $A\approx 1.34$ (a 34% brightening).
@@
Hence, for Earth-mass lenses such as small rogue planets ($M\sim M_{\oplus}$), this is on the order of $0.1\,\mathrm{ns}$, while for black holes such as those found in the LIGO band Nitz et al. (2021); Abbott et al. (2023) ($M\sim 30M_{\odot}$), this is on the order of $1\,\mathrm{ms}$, evaluated at $u\approx 1$.
@@ II.2 Problem setup for delay finding
Note that, in reality, $t_{0}$ is as a uniformly random quantity because one can never predict at what time a photon arrives. Therefore, the most rigorous way of expressing the state is a density operator $\rho(\Delta t)$, defined as a (classical) uniform mixture over all $|\ket{\phi(t_{0},\Delta t)}\bra{\phi(t_{0},\Delta t)}$ with fixed $\Delta t$ and $t_{0}\in[0,T]$, i.e.,
@@ III Frequency-domain interference
In this section, we propose our quantum-inspired algorithm for sample-efficient delay finding. We first introduce the main theoretical intuition in Section III.1 and Section III.2, using the fact that a fixed delay between two signals in the time domain corresponds to a modulation in the frequency domain. Next, we describe our algorithm based on frequency-basis measurements in Section III.3. Then, in the context of our Fourier-basis analysis, we explain in Section III.4 how the variance in $\Delta t$ caused by the finite-source effect may destroy the signal from an information-theoretic perspective. Finally, in Section III.5, we discuss the performance of our algorithm in the presence of noise.
@@ III.1 Fully classical picture
We observe that $|F(\tau)|^{2}$ has peaks at $\tau\rightarrow 0$ and $\tau\rightarrow\Delta t$. Therefore, the time delay can be directly read out from the Fourier transform of the power spectrum as a nonzero peak. This gives a straightforward approach to measuring the time delay by measuring the whole power spectrum of the optical signal, as in Refs. Deguchi and Watson (1986); Peterson and Falk (1991); Gould (1992); Ulmer and Goodman (1995); Katz et al. (2018); Eichler (2017); Jow et al. (2020); Wucknitz et al. (2021).
@@
Before moving on, we emphasize that the modulation $\cos(\omega\Delta t)$ in the power spectrum has an extremely short period, $1/\Delta t$. (Note that this “period” has units of $\mathrm{time}^{-1}$ because the oscillation is in the frequency domain.) For typical microlenses, $\Delta t$ can be as high as $10^{-3}\,\mathrm{s}$, corresponding to a $1\,\mathrm{kHz}$ period. Therefore, according to the Nyquist–Shannon sampling theorem, one must be able to obtain frequency information with up to kHz-level error to find $\Delta t$, regardless of how the data are processed. Furthermore, due to the uncertainty principle, measuring the frequency with precision $O(1/\Delta t)$ implies that a measurement of the time has error $\Omega(\Delta t)$. This is necessary for a lensed photon with $\Delta t$ time delay, and implies that the device must not be able to tell the difference between two time points in the same $\Delta t$ interval (see e.g. Ref. Vittorini et al. (2014) for a similar effect with the roles of time and frequency domains reversed). In other words, a device with $1/\Delta t$ frequency resolution must somehow “store” a photon for time at least $\Delta t$.
@@ III.2 Photonic (quantum) picture
In addition, we can show that the density operator $\rho(\Delta t)$ is diagonal in the Fourier basis for large $T$. (In fact, it must be diagonal for large $T$ in the Fourier basis because $\rho(\Delta t)$ is time-translation invariant.) To see this, we simply express the density operator in the frequency domain:
$\quad\rho(\Delta t)$ (37)
@@
$\quad\quad\int_{0}^{T}p(t_{0})e^{\mathrm{i}(\omega_{1}-\omega_{2})t_{0}}dt_{0}.$
@@
$\int_{0}^{T}p(t_{0})e^{i(\omega_{1}-\omega_{2})t_{0}}dt_{0}$ $=\frac{e^{i(\omega_{1}-\omega_{2})}-1}{iT(\omega_{1}-\omega_{2})}$ (38)
$=O\mathopen{}\mathclose{{\left(\frac{1}{T}}}\right)\rightarrow 0.$
Therefore,
$\quad\rho(\Delta t)$ (39)
@@ III.3 The sample-efficient algorithm
Recall that the value of $\Delta t$ can be directly read from a peak in the Fourier transform of the power spectrum, as shown in Eqs. (32, 36). Our sample-efficient algorithm reconstructs the peak from a limited number of frequency-domain measurement outcomes. We are inspired by maximum likelihood estimation algorithm and studies of the dihedral hidden subgroup problem Ettinger and Høyer (2000) to propose the score function
@@ III.4 The finite-source effect
Now, compared with $p(\omega|\Delta t)$ in Eq. (35), the $\cos(\omega\Delta t_{0})$ oscillation, which carries the information about $\Delta t_{0}$, is exponentially suppressed in the marginal distribution when $\delta_{\Delta t}\gtrsim 1/\omega_{0}$. Note that $1/\omega_{0}$ can be extremely tiny—as an example, $1/\omega_{0}\sim 10^{-14}\,\mathrm{s}$ for visible light.
@@
$\quad F_{\Delta t_{0},\delta_{\Delta t}}(\tau)$ (48)
$=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}p(\omega)e^{i\omega\tau}d\omega$
@@ III.6 Unequal magnification
$=\frac{1}{\sqrt{A}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{\left[\sqrt{A_{+}}\alpha(t-t_{0})e^{-i\omega_{0}t}+}}\right$.
@@ IV Sample complexity lower bound
Using $\ln(AB)=\ln A+\ln B$ and $\int_{0}^{T}\cos(\nu\Delta t)d\Delta t$ is much less than $T$ when $T\gg 2\pi/\nu$, the above integration can be simplified as
$\quad S_{\mathrm{right}}$ (64)
$\approx S_{\mathrm{left}}-\frac{1}{T}\int_{0}^{T}d\Delta t\int_{-\infty}^{\infty}d\nu\frac{t_{c}}{\sqrt{\pi}}e^{-t_{c}^{2}(\omega_{0}-\nu)^{2}}\cdot$
@@ V Quantum undersampling
Given the condition that all wave packets are of $1/t_{c}$ width in the frequency domain, it suffices to only distinguish the $\sim\frac{1/t_{c}}{1/T}=T/t_{c}$ frequency modes within the wave packet if one can localize the photon’s frequency to a range of $1/t_{c}$ width. This localization process can be thought of as a non-demolition measurement of photon frequency with $1/t_{c}$ resolution. One possible way to realize such a measurement is via a two-step process, which first splits photons into ports with $1/t_{c}$ frequency range (using e.g. a diffraction grating) and then uses non-demolition photon detectors to determine which port the photon is in. Once the photon is localized to a frequency range with $1/t_{c}$ resolution, it can be sent to a single-photon spectrometer with $1/t_{c}$ bandwidth and $1/T$ resolution. One spectrometer suffices provided that a coherent frequency converter is used or the spectrometer has tunable frequency range. Alternatively, one can use $\sim\frac{\omega_{\mathrm{max}}}{1/t_{c}}=\omega_{\mathrm{max}}t_{c}$ spectrometers. Combining the photon splitting device with spectrometry still distinguishes $\sim T\omega_{\mathrm{max}}$ modes, although the implementation is potentially less challenging than direct frequency measurement with a single broadband high-resolution single-photon spectrometer.
In this section, we present another way—that doesn’t use standard high-resolution spectrometers—of implementing the delay-finding procedure under the condition that we have already performed non-demolition measurement of the photon’s frequency with $\sim 1/t_{c}$ resolution. In particular, we propose to then use quantum information processing techniques to store the discretized and undersampled wave function of a photon in the time domain in a quantum memory (Section V.1) and perform the quantum Fourier transform on it (Section V.2).

### v2

@@ II Preliminaries
Throughout this paper, we adopt a few non-SI units that are standard in astronomy. Specifically, we use $1\,\mathrm{pc}=3.0857\times 10^{16}\,\mathrm{m}$ and $1\,\mathrm{erg}=10^{-7}\,\mathrm{J}$.
@@ II.1 Gravitational lensing time delays
When light emitted by a distant source passes near a large distribution of mass on its way to Earth, the path of the light is altered in an effect known as gravitational lensing. For sufficiently large masses, the distortion is large enough to form multiple images of the same source on the sky. However, at lower masses, these images cannot be individually resolved—they overlap with each other and with the true position of the source, causing the source to appear brighter. This transient magnification of a source is known as gravitational microlensing [40]. Such microlensing events provide one of the few ways to detect non-luminous astrophysical bodies.
For the purposes of this paper, we will restrict ourselves to a fiducial example, namely a star in the Galactic Bulge of the Milky Way as the source and a dark, isolated object such as a black hole as the lens. This example is of particular interest, as the Galactic Bulge has been the target of decades of ground-based microlensing searches [41, 42, 43], and such surveys have yielded significant discoveries of dark objects such as isolated black holes [44, 15, 16] and free-floating planets [45, 14, 46, 47, 48, 49, 50, 51, 52, 13, 53, 54]. In the next few years, future missions such as NASA’s Nancy Grace Roman Space Telescope (henceforth, Roman) [55, 56] and the Chinese National Space Agency’s Earth 2.0 satellite [57] will conduct the first ever dedicated space-based microlensing surveys. They are expected to discover orders of magnitude more dark astrophysical bodies than current ground-based surveys [56, 58].
However, as powerful a tool as microlensing is for discovering dark astrophysical objects, it suffers from inherent degeneracies that make measuring the underlying properties of the lens, such as its mass, quite challenging [59]. This can be seen from the fact that the primary observable associated with microlensing is a quantity called the Einstein crossing time that corresponds to the approximate duration of the lensing event. The Einstein crossing time is defined as the time for the source to cross the angular Einstein radius $\theta_{E}$, which is the region of the sky surrounding the lens in which the source is magnified by the lens. The Einstein crossing time $t_{E}$ therefore depends on the distance to the lens ($D_{L}$), the distance to the source ($D_{S}$), the mass of the lens ($M$), and the relative proper motion of the lens and source $\mu_{\text{rel}}$:
@@
The impact parameter can be independently measured from the magnification curve, as the magnification in the point-source point-lens regime is [11]
@@
where $u$ varies with time as the source location approaches its minimum impact parameter $u_{0}$ and then moves away from the lens axis. We see that if $u_{0}=0$, the magnification is infinite, as expected [40]. Given the magnification curve and the time delay, it is clear that one can solve for $M$, the lens mass, directly, breaking the inherent degeneracy present in the light curve alone.
The behavior of Eqs. (7) and (8), as plotted in Fig. 2, constrains the regime in which microlensing is useful. For a microlensing event to be detectable, we normally require $u\approx 1$ (though with future space-based missions, this may be relaxed). Additionally, note that Eq. (8) is the total magnification, i.e., the sum of the major and minor image contributions given by $A_{\pm}(u)=\frac{1}{2}A(u)\pm\frac{1}{2}$. The analysis is simpler when there is comparable flux from the two images, i.e., $u\approx 1$. For both of these reasons, we mainly consider examples in which $u\approx 1$, hence $\Delta t\approx\frac{8GM}{c^{3}}$ and $A\approx 1.34$ (a $34\,\%$ brightening).
@@
Hence, for Earth-mass lenses such as small rogue planets ($M\sim M_{\oplus}$), this is on the order of $0.1\,\mathrm{ns}$, while for black holes such as those found in the LIGO band [60, 61] ($M\sim 30M_{\odot}$), this is on the order of $1\,\mathrm{ms}$, evaluated at $u\approx 1$.
@@ II.2 Problem setup for delay finding
Note that, in reality, $t_{0}$ is as a uniformly random quantity because one can never predict at what time a photon arrives. Therefore, the most rigorous way of expressing the state is a density operator $\rho(\Delta t)$, defined as a (classical) uniform mixture over all $|\ket{\phi(t_{0},\Delta t)}\bra{\phi(t_{0},\Delta t)}$ with fixed $\Delta t$ and $t_{0}\in[0,T_{\mathrm{w}}]$ (with $T_{\mathrm{w}}\gg\Delta t$), i.e.,
@@ III Frequency-domain interference
In this section, we propose our quantum-inspired algorithm for sample-efficient delay finding. We first introduce the main theoretical intuition in Section III.1 and Section III.2, using the fact that a fixed delay between two signals in the time domain corresponds to a modulation in the frequency domain. Next, we describe our algorithm based on frequency-basis measurements in Section III.3. Then, in the context of our Fourier-basis analysis, we explain in Section III.4 how the variance in $\Delta t$ caused by the finite-source effect may destroy the signal from an information-theoretic perspective. Next, in Section III.5 and Section III.6, we discuss the performance of our algorithm in the presence of noise and unequal magnification, respectively. Finally, we discuss the realistic broadband scenario where photons have different carrier frequencies in Section III.7.
@@ III.1 Fully classical picture
We observe that $|F(\tau)|^{2}$ has peaks at $\tau\rightarrow 0$ and $\tau\rightarrow\Delta t$. Therefore, the time delay can be directly read out from the Fourier transform of the power spectrum as a nonzero peak. This gives a straightforward approach to measuring the time delay by measuring the whole power spectrum of the optical signal, as in Refs. [22, 23, 24, 25, 26, 27, 28, 29].
@@
Before moving on, we emphasize that the modulation $\cos(\omega\Delta t)$ in the power spectrum has an extremely short period, $1/\Delta t$. (Note that this “period” has units of $\mathrm{time}^{-1}$ because the oscillation is in the frequency domain.) For typical microlenses, $\Delta t$ can be as high as $10^{-3}\,\mathrm{s}$, corresponding to a $1\,\mathrm{kHz}$ period. Therefore, according to the Nyquist–Shannon sampling theorem, one must be able to obtain frequency information with up to kHz-level error to find $\Delta t$, regardless of how the data are processed. Furthermore, due to the uncertainty principle, measuring the frequency with precision $O(1/\Delta t)$ implies that a measurement of the time has error $\Omega(\Delta t)$. This is necessary for a lensed photon with $\Delta t$ time delay, and implies that the device must not be able to tell the difference between two time points in the same $\Delta t$ interval (see e.g. Ref. [62] for a similar effect with the roles of time and frequency domains reversed). In other words, a device with $1/\Delta t$ frequency resolution must somehow “store” a photon for time at least $\Delta t$.
@@ III.2 Photonic (quantum) picture
In addition, we can show that the density operator $\rho(\Delta t)$ is diagonal in the Fourier basis for large time window $T_{\mathrm{w}}$. (In fact, it must be diagonal for large $T$ in the Fourier basis because $\rho(\Delta t)$ is time-translation invariant.) To see this, we simply express the density operator in the frequency domain:
@@
$\quad\quad\int_{0}^{T_{\mathrm{w}}}p(t_{0})e^{\mathrm{i}(\omega_{1}-\omega_{2})t_{0}}dt_{0}.$
@@
$\int_{0}^{T_{\mathrm{w}}}p(t_{0})e^{i(\omega_{1}-\omega_{2})t_{0}}dt_{0}$ $=\frac{e^{i(\omega_{1}-\omega_{2})T_{\mathrm{w}}}-1}{iT_{\mathrm{w}}(\omega_{1}-\omega_{2})}$ (38)
$=O\mathopen{}\mathclose{{\left(\frac{1}{T_{\mathrm{w}}}}}\right)\rightarrow 0.$
Therefore, when $T_{\mathrm{w}}$ is sufficiently large,
$\quad\rho(\Delta t)\approx\rho_{\mathrm{diag}}(\Delta t)$ (39)
@@ III.3 The sample-efficient algorithm
Recall that the value of $\Delta t$ can be directly read from a peak in the Fourier transform of the power spectrum, as shown in Eqs. (32, 36). Our sample-efficient algorithm reconstructs the peak from a limited number of frequency-domain measurement outcomes. We are inspired by maximum likelihood estimation algorithm and studies of the dihedral hidden subgroup problem [33] to propose the score function
@@ III.4 The finite-source effect
Now, compared with $p(\omega|\Delta t)$ in Eq. (35), the $\cos(\omega\Delta t_{0})$ oscillation, which carries the information about $\Delta t_{0}$, is exponentially suppressed in the marginal distribution when $\delta_{\Delta t}\gtrsim 1/\omega_{0}$. Note that $1/\omega_{0}$ can be extremely tiny—as an example, $1/\omega_{0}\sim 10^{-15}\,\mathrm{s}$ for visible light.
@@
$F_{\Delta t_{0},\delta_{\Delta t}}(\tau)$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}p(\omega)e^{i\omega\tau}d\omega$ (48)
@@ III.6 Unequal magnification
III.7 Broadband input and coherence time
In the previous analysis, we assume that all photons are within the frequency range set by the bandwidth of the single-photon spectrometer. In other words, we assume all photons are Gaussian wave packets centered at $\omega_{0}$ with width $\sim 1/t_{c}$. However, as later discussed in Section VII, in a realistic setting, a typical photon source emits broadband light of bandwidth $\sim 10^{14}\,\mathrm{Hz}$, and we only have access to several hundred photons that come from this bandwidth. In this case, the carrier frequency $\omega_{0}$ can be drastically different for different photons, and we must consider the potential effect of these broadband input photons.
Fortunately, Eq. (39) indicates that the $1+\cos(\omega\Delta t)$ modulation in the wave function of the incoming photon is independent of the carrier frequency $\omega_{0}$. Therefore, the density matrix of the broadband input photons can be written as the integration over $\rho(\Delta t)$ of all possible carrier frequencies, i.e.,
[…]
where $\mathcal{A}$ is the normalization factor, $\omega_{\mathrm{L}},\omega_{\mathrm{R}}$ are the lower and upper limits of the passband, and $p_{\mathrm{B}}(\omega)$ is the envelope of the frequency distribution, which is usually the black-body spectrum and can be approximately considered as a constant. With the constant $p_{\mathrm{B}}$ assumption, the classical communication channel can be written as
$p(\omega|\Delta t)\approx\frac{1+\cos(\omega\Delta t)}{\omega_{\mathrm{R}}-\omega_{\mathrm{L}}}.$ (58)
Now, we observe that inferring $\Delta t$ using photons with carrier frequency sampled from a wide range is equivalent to using photons with a broadband wave packet profile. This implies that the coherence time in our broadband time-delay estimation task is $\sim 1/(\omega_{\mathrm{R}}-\omega_{\mathrm{L}})\sim 10^{-15}\,\mathrm{s}$, independent of the bandwidth of every single-photon spectrometer. Indeed, one can compute the expectation value of the score function with $p_{\mathrm{B}}$:
$\mathbb{E}$ $[\cos(\nu_{j}\tau)]=\int_{\omega_{\mathrm{L}}}^{\omega_{\mathrm{R}}}\cos(\nu_{j}\tau)p_{\mathrm{B}}(\omega|\Delta t)d\nu_{j}$ (59)
$=\frac{\mathopen{}\mathclose{{\left[\sin(\omega(\tau-\Delta t))}}\right]_{\omega_{\mathrm{L}}}^{\omega_{\mathrm{R}}}}{2(\omega_{\mathrm{R}}-\omega_{\mathrm{L}})(\tau-\Delta t)}+O\mathopen{}\mathclose{{\left(((\omega_{\mathrm{R}}-\omega_{\mathrm{L}})\tau)^{-1}}}\right),$
whose absolute value is close to 0 when $(\omega_{\mathrm{R}}-\omega_{\mathrm{L}})|\tau-\Delta t|\gg 1$ and is close to $1/2$ when $(\omega_{\mathrm{R}}-\omega_{\mathrm{L}})|\tau-\Delta t|\leq 1$. Therefore, the precision of $\Delta t$ estimation is set by the coherence time $t_{c}\sim 1/(\omega_{\mathrm{R}}-\omega_{\mathrm{L}})\sim 10^{-15}\,\mathrm{s}$. Henceforth in this paper, when considering the broadband input scenario, $t_{c}$ is the coherence time determined by the signal bandwidth. Since the former is not tunable for a given passband, the number of required photons in the broadband case (which scales with $\log(T/t_{c})$) is solely determined by $T$, the upper limit of $\Delta t$.
@@ IV Sample complexity lower bound
Using $\ln(AB)=\ln A+\ln B$ and the fact that $\int_{0}^{T}\cos(\nu\Delta t)d\Delta t$ is much less than $T$ when $T\gg 2\pi/\nu$, the above integration can be simplified as
$S_{\mathrm{right}}$ $\approx S_{\mathrm{left}}-\frac{1}{T}\int_{0}^{T}d\Delta t\int_{-\infty}^{\infty}d\nu\frac{t_{c}}{\sqrt{\pi}}e^{-t_{c}^{2}(\omega_{0}-\nu)^{2}}\cdot$ (67)
@@ V Quantum undersampling
However, suppose all wave packets are of $1/t^{\prime}_{c}$ width (with $t^{\prime}_{c}\gg t_{c}$) in the frequency domain, it suffices to only distinguish the $\sim\frac{1/t^{\prime}_{c}}{1/T}=T/t^{\prime}_{c}$ frequency modes within the wave packet if one can localize the photon’s frequency to a range of $1/t^{\prime}_{c}$ width. This localization process can be thought of as a non-demolition measurement of photon frequency with $1/t^{\prime}_{c}$ resolution. One possible way to realize such a measurement is via a two-step process, which first splits photons into ports with $1/t^{\prime}_{c}$ frequency range (using e.g. a diffraction grating) and then uses non-demolition photon detectors to determine which port the photon is in. Once the photon is localized to a frequency range with $1/t^{\prime}_{c}$ resolution, it can be sent to a single-photon spectrometer with $1/t^{\prime}_{c}$ bandwidth and $1/T$ resolution. One spectrometer suffices provided that a coherent frequency converter is used or the spectrometer has tunable frequency range. Alternatively, one can use $\sim\frac{\omega_{\mathrm{max}}}{1/t^{\prime}_{c}}=\omega_{\mathrm{max}}t^{\prime}_{c}$ spectrometers. Combining the photon splitting device with spectrometry still distinguishes $\sim T\omega_{\mathrm{max}}$ modes, although the implementation is potentially less challenging than direct frequency measurement with a single broadband high-resolution single-photon spectrometer.
In this section, we present another way—that doesn’t use standard high-resolution spectrometers—of implementing the delay-finding procedure under the condition that we have already performed non-demolition measurement of the photon’s frequency with $\sim 1/t^{\prime}_{c}$ resolution. In particular, we propose to then use quantum information processing techniques to store the discretized and undersampled wave function of a photon in the time domain in a quantum memory (Section V.1) and perform the quantum Fourier transform on it (Section V.2).